# Yibo Shi, Braghadeesh Lakshminarayanan and Cristian R. Rojas

### v1

@@ IV Numerical Illustrations
    NLSS system: a sampling example where the FIM is estimated via the particle filter method, showing that Algorithm 1 remains effective with estimated $\boldsymbol{J}_{\boldsymbol{\theta}}$;
@@
    An application of parameter estimation: a practical setting in which Jeffreys prior sampling outperforms a uniform prior for synthetic-data generation, improving parameter estimation performance.
We now detail each example in turn.
@@ IV-A1 Experiment setup
In this experiment, we apply Algorithm 1 to sample from Jeffreys prior for the Coin-Bending Model introduced in [25]. In this model, the probability $q(\varphi)$ of obtaining heads depends on the bending angle $\varphi$ according to
@@
Each coin toss yields a binary outcome random variable $Y$ where
$Y=\left\{\begin{array}[]{l}1,\text{ with probability }\frac{1}{2}+\frac{1}{2}\left(\frac{\varphi}{\pi}\right)^{3},\\ 0,\text{ with probability }\frac{1}{2}-\frac{1}{2}\left(\frac{\varphi}{\pi}\right)^{3},\end{array}\right.$
with 1 denoting heads. Therefore, we can derive the score function in closed form:
@@
Assuming consistency as $n\rightarrow\infty$, we treat $\pi(\varphi)\propto\sqrt{\hat{J}_{\varphi}}$ as the ground truth, we define the potential $V(\varphi)=-\frac{1}{2}\ln\hat{J}_{\varphi}$ with the gradient given explicitly by
@@
[…]
@@ IV-A2 Simulation results
The simulation results are shown in Fig. 1. As illustrated, the empirical distribution of samples obtained by our algorithm gives better match of the theoretical Jeffreys prior as the number of samples increases (from Row 1 to Row 3). Moreover, even with a relatively small sample size (e.g., 1,000 samples), the empirical distribution already closely resembles the exact Jeffreys prior, indicating rapid convergence. Additionally, the difference in the sampled realizations (100 sampled points) reveals the stochastic nature of the chain and demonstrates good mixing behavior, reflecting adequate exploration of the parameter space.
This example provides a clear validation of our proposed MALA-based sampler, as it allows direct comparison between the empirical and exact prior distributions. Furthermore, this one-dimensional scenario enables straightforward assessment of mixing quality and acceptance rates, establishing a solid baseline for applying our approach in more complex, high-dimensional settings.
@@ IV-B Sampling Jeffreys Prior for a Dynamical System
Figure 2: Histogram of 10000 samples of Jeffreys prior and the shape of the true Jeffreys prior (red curve) from the SV Model.
@@ IV-B1 Experiment setup
In this example, we demonstrate the performance of Algorithm 1 by sampling from the Jeffreys prior for an NLSS model, where the FIM is numerically approximated using a particle filter. Specifically, we consider the Hull–White stochastic volatility (SV) model [26, 16] defined as:
$x_{t+1}\mid x_{t}$ $\sim\mathcal{N}\left(\varphi x_{t}+\rho u_{t},\sigma_{v}^{2}\right),$
$y_{t}\mid x_{t}$ $\sim\mathcal{N}\left(0,\beta^{2}\exp\left(x_{t}\right)\right),$
where the control input $u_{t}$ is generated by samples from a standard normal distribution. Our goal is to sample the Jeffreys prior $\pi(\varphi)$ for the system parameter $\varphi$ corresponding to the system dynamics, constrained to the feasible region $\Phi_{c}=[0.3,0.9]$, ensuring system stability. For simulation, the rest of model parameters are set as
[…]
and the system is simulated for $T=1000$ time steps.
Since the Fisher information $J(\varphi)$ lacks an analytical form for this model, we estimate it using a particle-filter–based method from [17], as described in Section III, employing $N_{p}=1000$ particles and averaging multiple Monte Carlo runs to obtain a reliable estimation $\hat{J}(\varphi)$. Accordingly, we compute the estimated shape of Jeffreys prior $\pi(\varphi)\propto\sqrt{\hat{J}_{\varphi}}$ and compare it with the distribution of the generated samples.
We define the corresponding potential function as $V(\varphi)=-\frac{1}{2}\ln\hat{J}(\varphi)$ and approximate its gradient $\widehat{\nabla V}(\varphi)$ using (12) and the one-point finite-difference estimator in (13).
@@ IV-B2 Simulation results
The histogram of the $N=10000$ generated samples is shown in Fig. 2, overlaid with the estimated shape of Jeffreys prior $\pi(\varphi)\propto\sqrt{\hat{J}_{\varphi}}$ computed from the particle filter estimation.
As expected, the estimated Jeffreys prior assigns higher probability near the boundary $\varphi\approx 0.9$, indicating greater parameter sensitivity, and lower probability near the boundary ($\varphi\approx 0.3$.
At the lower boundary $\varphi=0.3$, the influence of $x_{t}$ on the evolution of the state is relatively small; the dynamics are then dominated by the input $u_{t}$ and the process noise $v_{t}$, making the likelihood less sensitive to changes in $\varphi$. In contrast, around $\varphi=0.9$, a small perturbation in $\varphi$ produces significant changes in the predicted states and, hence, in the likelihood function. This leads to a higher Fisher information in that region. Thus, the estimated prior shape confirms that the particle-filter-based FIM estimation is consistent with theoretical expectations.
Moreover, the close agreement observed in Fig. 2 between the histogram of generated samples and the estimated Jeffreys prior verifies that our proposed algorithm reliably samples from the Jeffreys prior even when the FIM is numerically approximated.
Overall, these results demonstrate the validity of our sampling framework and highlight the advantage of leveraging the geometry encoded by the Jeffreys prior for parameter inference and experimental design.
@@ IV-C An Application of Jeffreys Prior to Parameter Estimation
Figure 3: (a) Heatmap of 1000 samples from Jeffreys prior distribution of the scale ($\eta$) and shape ($\gamma$) parameters from the Weibull distribution model; (b) Scatter plot of estimated scale parameter $\hat{\eta}$ based on the uniform prior and Jeffreys prior vs. its true value; (c) Scatter plot of estimated shape parameter $\hat{\gamma}$ based on the uniform prior and Jeffreys prior vs. its true value. The red dashed line corresponds to an oracle estimate, which knows the true value of the parameter.
@@ IV-C1 Experiment setup
In this example, we illustrate the advantages of sampling from Jeffreys’ prior within the Two-Stage (TS) estimation framework, using parameter estimation for a Weibull distribution as a test case. TS provides an effective validation scenario, although the framework itself is well-established. In TS, given a parametric model $p(\cdot;\boldsymbol{\theta})$ with parameters $\boldsymbol{\theta}\in\Theta\subseteq\mathbb{R}^{d}$, one constructs a synthetic training dataset of parameter-data pairs
$\{(\boldsymbol{\theta}_{i},\boldsymbol{y}_{i})\}_{i=1}^{M_{\theta}},$
where data samples $\boldsymbol{y}_{i}$ are generated from $p(\cdot;\boldsymbol{\theta}_{i})$. A supervised learning method then builds an estimator:
$\hat{\boldsymbol{\theta}}(\boldsymbol{y})=\boldsymbol{g}(\boldsymbol{h}(\boldsymbol{y})),$
with a fixed compression function $\boldsymbol{h}$ and a regression-based function $\boldsymbol{g}$. For additional details on the implementation setup of TS, we refer the reader to [27].
In the context of TS, the distribution of the synthetic parameter samples is crucial. We advocate sampling $\boldsymbol{\theta}$ according to Jeffreys prior, $\pi(\boldsymbol{\theta})$ naturally reflects the local identifiability structure of the model. In contrast, a uniform prior may not account for how informative different regions of the parameter space are, potentially leading to suboptimal training datasets.
The advantage of our approach is that, by sampling the synthetic $\boldsymbol{\theta}_{i}$ values from Jeffreys prior using our proposed algorithm, the resulting training dataset better represents regions where the data is most informative about the parameters.
For the numerical example, we consider the Weibull distribution, widely used in reliability engineering, whose probability density function is given by
@@
where $\eta>0$ is the scale parameter and $\gamma>0$ is the shape parameter. Thus, the parameter vector is $\boldsymbol{\theta}=[\eta,\gamma]^{T}$.
@@ IV-C2 Simulation results
We validate the TS estimators trained under uniform and Jeffreys priors using a validation set consisting of $1000$ parameter points $\boldsymbol{\theta}_{\ell}=[\eta_{\ell},\gamma_{\ell}]^{T}$ uniformly sampled from $[1,20]\times[1,20]$. For each $\boldsymbol{\theta}_{\ell}$ synthetic data $\{y^{i}_{\ell}\}_{i=1}^{M}$ generated from the Weibull model. We evaluate two classes of TS estimators based on the samples from the uniform and Jeffreys priors respectively.
Fig. 3(a) shows the sampled Jeffreys prior distribution in the $(\gamma,\eta)$ parameter space. Notably, Jeffreys prior emphasizes lower values of $\gamma$ while remaining relatively uniform along $\eta$, aligning well with the Weibull model’s FIM structure, which indicates higher sensitivity (information content) at smaller $\gamma$.
@@
However, notable differences arise for the shape parameter $\gamma$, as seen in Fig. 3(c). Near low values of $\gamma<5$, the uniform-based estimator exhibits significant variance and bias. In contrast, the Jeffreys-based estimator demonstrates considerably improved accuracy and precision in this region, owing to the fact that Jeffreys prior focuses sampling efforts where the Fisher information is greatest. Remarkably, despite allocating fewer training samples to higher $\gamma$ regions, the Jeffreys-based estimator maintains effective generalization in those regions, underscoring the advantage of information-driven weighting in the parameter space.
Overall, these experiments demonstrate that sampling according to Jeffreys prior within the TS framework substantially enhances estimation accuracy and robustness for shape-dominated parameters, while maintaining comparable performance to the uniform-based estimator along directions of weaker sensitivity. Therefore, the application of Jeffreys prior not only leads to better TS estimators for the Weibull distribution but also suggests that our approach may have broader applicability in synthetic data generation and parameter estimation for complex models.
@@ V Conclusions
In this paper, we have proposed a MALA-based sampling scheme designed to generate samples from Jeffreys prior. Our approach involves imposing Jeffreys prior as the stationary distribution of a Langevin-based MCMC, where each update step of the Markov chain is determined by the gradient of the logarithm of the determinant of the Fisher Information matrix. Furthermore, for the case when this matrix cannot be computed analytically, we have employed a particle filter algorithm to approximate the score function, thereby estimating the gradient of the log determinant of the FIM. We have validated our sampling scheme through several numerical examples, including one that has demonstrated how the diversity inherent in Jeffreys prior can enhance the estimates produced by a data-driven estimator. In future work, we plan to further explore the use of this sampling scheme for parameter estimation in dynamical systems, particularly using data-driven estimators that require minimal training data, by leveraging the diversity of samples promoted by Jeffreys prior.

### v3

@@ IV Numerical Illustrations
    NLSS system: a sampling example where the FIM is estimated via the PF method, showing that Algorithm 1 remains effective with estimated $\boldsymbol{J}_{\boldsymbol{\theta}}$;
@@
    Parameter estimation: a practical setting in which Jeffreys prior sampling outperforms a uniform prior in data-driven parameter estimation performance.
@@ IV-A1 Experiment setup
In this experiment, we apply Algorithm 1 to sample from Jeffreys prior for a Coin-Bending Model [23], where the probability $q(\varphi)$ of obtaining heads depends on the bending angle $\varphi$ according to
@@
Each coin toss yields a binary random variable $Y$, where $Y=1$ with probability $\frac{1}{2}+\frac{1}{2}\left(\frac{\varphi}{\pi}\right)^{3}$, and $Y=0$ otherwise, with 1 denoting heads. Its score function is
@@
Assuming consistency as $n\rightarrow\infty$, we treat $\pi(\varphi)\propto\sqrt{\hat{J}_{\varphi}}$ as the ground truth, and define the potential $V(\varphi)=-\frac{1}{2}\ln\hat{J}_{\varphi}$ whose gradient given explicitly by
@@
We run Algorithm 1 to generate samples from $\pi(\varphi)$ within the constrained set $\varphi\in\Phi_{c}=[2,3]$. Three experiments are performed with $N\in\{100,\ 1000,\ 10000\}$ samples, each repeated for 10 independent realizations. The resulting empirical distributions are then compared with the estimated Jeffreys prior to verify convergence and consistency.
@@ IV-A2 Simulation results
The simulation results are shown in Fig. 1, which shows that the empirical distribution of samples obtained by our algorithm matches the theoretical Jeffreys prior as the number of samples increases (from Row 1 to Row 3). Moreover, only a relatively small sample size (e.g., 1,000 samples) yields a very good match to the exact Jeffreys prior, indicating rapid convergence. Additionally, the difference in the sampled realizations (for $N=100$) reveals the stochastic nature of the chain, demonstrating good mixing behavior and adequate exploration of the parameter space.
This example provides a clear validation of our proposed MALA-based sampler, as it allows direct comparison between the empirical and exact prior distributions. Furthermore, this one-dimensional scenario enables straightforward assessment of mixing quality and acceptance rates.
@@ IV-B1 Experiment setup
In this example, we apply Algorithm 1 on an NLSS model, where the FIM is numerically approximated using a PF. Specifically, we consider the Hull–White stochastic volatility (SV) model [24, 16]
$x_{t+1}\mid x_{t}\sim\mathcal{N}\left(\varphi x_{t}+\rho u_{t},\sigma_{v}^{2}\right),\quad y_{t}\mid x_{t}\sim\mathcal{N}\left(0,\beta^{2}e^{x_{t}}\right),$
where the control input $u_{t}$ is white noise of standard normal distribution. Our goal is to sample the Jeffreys prior $\pi(\varphi)$ for the system parameter $\varphi$ corresponding to the system dynamics, constrained to the feasible region $\Phi_{c}=[0.3,0.9]$, ensuring system stability. For simulation, the remaining model parameters are set as $\rho,\ \sigma_{v},\ \beta]=[0.2,\ 0.5,\ 0.7]$, and the system is simulated for $T=1000$ time steps.
Since the FIM $J(\varphi)$ lacks an analytical form for this model, we estimate it using a PF–based method from [17], as described in Section III, employing $N_{p}=1000$ particles and averaging multiple Monte Carlo runs to obtain a reliable estimate $\hat{J}(\varphi)$. Accordingly, we compute the estimated shape of Jeffreys prior $\pi(\varphi)$ and compare it with the distribution of the generated samples.
We define the potential function as $V(\varphi)=-\frac{1}{2}\ln\hat{J}(\varphi)$ and approximate its gradient $\widehat{\nabla V}(\varphi)$ using (12) and (13).
@@ IV-B2 Simulation results
A histogram of the $10000$ generated samples is shown in Fig. 2, overlaid with the estimated shape of Jeffreys prior $\pi(\varphi)$ computed from the PF estimation.
As expected, the estimated Jeffreys prior assigns higher probability near the boundary $\varphi=0.9$, indicating greater parameter sensitivity, and lower probability near $\varphi=0.3$.
Around $\varphi=0.3$, the influence of $x_{t}$ on the evolution of the state is relatively small; the dynamics are then dominated by the input $u_{t}$ and the process noise $v_{t}$, making the likelihood less sensitive to changes in $\varphi$. In contrast, around $\varphi=0.9$, a small perturbation in $\varphi$ produces significant changes in the predicted states and, hence, in the likelihood function, leading to a higher FIM. Thus, the PF-based FIM estimation is consistent with theoretical expectations.
Moreover, the close agreement observed in Fig. 2 between the histogram of generated samples and the estimated Jeffreys prior verifies that our proposed algorithm reliably samples from Jeffreys prior even when the FIM is numerically approximated. Overall, the results from experiments 1 and 2 demonstrate the validity of our sampling framework.
Figure 2: Histogram of 10000 samples of Jeffreys prior and the shape of the PF-estimated Jeffreys prior (red curve) from the SV Model.
@@ IV-C An Application of Jeffreys Prior to Parameter Estimation
Figure 3: (a) Heatmap of 1000 samples from Jeffreys prior distribution of the scale ($\eta$) and shape ($\gamma$) parameters from a Weibull distribution; (b) Scatter plot of estimated scale parameter $\hat{\eta}$ based on the uniform prior and Jeffreys prior vs. its true value; (c) Scatter plot of estimated shape parameter $\hat{\gamma}$ based on the uniform prior and Jeffreys prior vs. its true value. The red dashed line corresponds to an oracle estimate, which knows the true value of the parameter.
@@ IV-C1 Experiment setup
In this example, we illustrate the advantages of sampling from Jeffreys prior within the Two-Stage (TS) estimation framework, using parameter estimation for a Weibull distribution as a test case. TS provides an effective validation scenario, although the framework itself is well-established. In TS, given a parametric model $p(\cdot;\boldsymbol{\theta})$ with parameters $\boldsymbol{\theta}\in\Theta\subseteq\mathbb{R}^{d}$, one constructs a synthetic training dataset of parameter-data pairs $\{(\boldsymbol{\theta}_{i},\boldsymbol{y}_{i})\}_{i=1}^{M_{\theta}}$, where data samples $\boldsymbol{y}_{i}$ are generated from $p(\cdot;\boldsymbol{\theta}_{i})$. A supervised learning method then builds an estimator of the form $\hat{\boldsymbol{\theta}}(\boldsymbol{y})=\boldsymbol{g}(\boldsymbol{h}(\boldsymbol{y}))$, with a fixed compression function $\boldsymbol{h}$ and a regression-based function $\boldsymbol{g}$. For additional details, see [25].
In the context of TS, the distribution of the synthetic parameter samples is crucial. We advocate sampling $\boldsymbol{\theta}$ according to Jeffreys prior, since $\pi(\boldsymbol{\theta})$ naturally reflects the local identifiability structure of the model, thus producing synthetic data which is most informative about the parameters. In contrast, a uniform prior may not account for how informative different regions of the parameter space are, potentially leading to suboptimal training datasets.
As a numerical example, we consider the Weibull distribution, widely used in reliability engineering, whose probability density function, parameterized by $\boldsymbol{\theta}=[\eta,\gamma]^{T}$, is given by
@@
where $\eta>0$ and $\gamma>0$ are the scale and shape parameters.
@@ IV-C2 Simulation results
We validate the TS estimators trained under uniform and Jeffreys priors using a validation set consisting of $1000$ parameter points $\boldsymbol{\theta}_{\ell}=[\eta_{\ell},\gamma_{\ell}]^{T}$ uniformly sampled from $1,20]\times[1,20]$. For each $\boldsymbol{\theta}_{\ell}$ synthetic data $\{y^{i}_{\ell}\}_{i=1}^{M}$ generated from the Weibull model. We evaluate two classes of TS estimators based on the samples from the uniform and Jeffreys priors respectively.
Fig. 3(a) shows the sampled Jeffreys distribution in the $(\gamma,\eta)$ parameter space. Notably, Jeffreys prior emphasizes lower values of $\gamma$ while remaining relatively uniform along $\eta$, aligning well with the Weibull model’s FIM structure, which indicates higher sensitivity at smaller $\gamma$.
@@
However, notable differences arise for the shape parameter $\gamma$, as seen in Fig. 3(c). For $\gamma<5$, the uniform-based estimator exhibits significant variance and bias. In contrast, the Jeffreys-based estimator has considerably higher accuracy in this region, since Jeffreys prior focuses sampling efforts where the FIM is greatest. Remarkably, despite allocating fewer training samples to higher $\gamma$ regions, the Jeffreys-based estimator maintains effective generalization in those regions, showing the advantage of information-driven weighting in the parameter space.
Overall, these experiments demonstrate that sampling according to Jeffreys prior within the TS framework substantially enhances estimation accuracy and robustness for shape-dominated parameters, while maintaining comparable performance to the uniform-based estimator along directions of weaker sensitivity. This also suggests that our approach may have broader applicability in synthetic data generation and parameter estimation for complex models.
@@ V Conclusions
In this paper, we have proposed a MALA-based scheme designed to generate samples from Jeffreys prior. Our approach imposes Jeffreys prior as the stationary distribution of a Langevin-based MCMC, where each update step of the Markov chain is determined by the gradient of the logarithm of the determinant of the FIM. Furthermore, for the case when this matrix cannot be computed analytically, we have employed a PF algorithm to approximate the score function, and thus the gradient of the log determinant of the FIM. We have validated our sampling scheme through several numerical examples, including one that has demonstrated how the diversity inherent in Jeffreys prior can enhance the performance of a data-driven estimator. In future work, we plan to further explore the use of this scheme for parameter estimation in dynamical systems, particularly using data-driven estimators that require minimal training data, by leveraging the sample diversity promoted by Jeffreys prior.